# VQE for Ising Model & A Comparative Analysis of Classical and Quantum Optimization Methods

### v1

@@ Abstract
In this study, we delved into several optimization methods, both classical and quantum, and analyzed the quantum advantage that each of these methods offered, and then we proposed a new combinatorial optimization scheme, deemed as QN-SPSA+PSR which combines calculating approximately Fubini-study metric (QN-SPSA) and the exact evaluation of gradient by Parameter-Shift Rule (PSR). The QN-SPSA+PSR method integrates the QN-SPSA computational efficiency with the precise gradient computation of the PSR, improving both stability and convergence speed while maintaining low computational consumption. Our results provide a new potential quantum supremacy in the VQE’s optimization subroutine and enhance viable paths toward efficient quantum simulations on Noisy Intermediate-Scale Quantum Computing (NISQ) devices. Additionally, we also conducted a detailed study of quantum circuit ansatz structures in order to find the one that would work best with the Ising model and NISQ, in which we utilized the symmetry of the investigated model.
@@
Ising Model, Variational Quantum Eigensolver, Quantum Optimization, Ansazt Construction
@@ I Introduction
From the early stage of studying quantum computers, that by its very quantum nature is promising for a bright new age of computation coming along with three crucial keys of quantum theory: quantum probabilistics, superposition, and entanglement. The distinct properties of quantum computers (QC) from classical computers make it be more unique in application, not only be an upgraded computational speed version of traditional computers. Currently, when the quantum hardware is rather limited on operational stats and noise resilience, despite of the ambiguity of the realization of such advantages, Variational Quantum Algorithms (VQAs) [1, 2] are thought to be the best at outperforming conventional computers, which are able to implement well on near-term quantum devices known as the so-called Noisy Intermediate-Scale Quantum (NISQ) computers.
@@
Nowadays, when the quantum revolution is being on track, big tech companies like IBM, Google, D-Wave, etc. compete to build their own quantum computers to yield quantum supremacy, where the first-ever experimental demonstration was achieved by the Google AI Quantum team [3], but that is still far from what we expect the quantum computer could do, while coherence time, connectivity of qubit, and qubit number limitations keep us from being able to successfully run a long depth circuit and produce a significant result on the present noisy device, in addition to quantum gate implementation issues. Those restrictions would yet be challenging to hardware scientists in the near future, posing obstacles that computer scientists would need to weigh against trainability, precision, and efficiency, nevertheless, these NISQ devices are capable of exploitation [4]. Standing out among quantum algorithms envisioned to beat classical computers, which are mostly designed for the fault-tolerant quantum computer, VQAs turns up regarded as an appropriate candidate compatible with the current defective quantum device to address these constraints [2]. The first two prominent applications of VQAs coming up are the Variational Quantum Eigensolver (VQE) and the Quantum Approximate Optimization Algorithm (QAOA). These inherit the core scheme of a VQA, which is to utilize the hybrid routine, leverage the versatility and computational power of classical computers to handle the computation processing, and use quantum devices to execute the quantum circuit. Hence, in this paper’s scope, we will travel through the standard procedures and investigate the properties of different optimization methods implemented in a VQE application.
@@
The structure of this paper is as follows, in Chapter 1, we provide an overview of the Variational Quantum Eigensolver (VQE) procedures. A brief introduction to the Transverse Ising Model and its symmetries is discussed in Chapter 2, which will be utilized to inform the ansatz structure. Subsequently, Chapter 3 delves into optimization methods, encompassing both classical and quantum approaches, where we will propose a new quantum optimization method. The numerical study and experiment are conducted in Chapter 4, and then we make the final conclusion in Chapter 5.
II Variational Quantum Eigensolver (VQE)
Figure 1: VQE Architecture.
@@ II Variational Quantum Eigensolver (VQE)
VQE was initially proposed by Peruzzo et al. [5, 6] as an efficient alternative way to compute the ground state energy of many quantum chemistry systems [7, 8] or many-body systems, i.e. ground-state He-H+ molecular energy, rather than Quantum Phase Estimation (QPE) [9] which demands impractically huge numbers of quantum gates. Consequently, VQE becomes a highly flexible, viable strategy with currently potential quantum resources that can outpace conventional computers. Following the success of running on the photonic quantum processor combined with the traditional devices, many research works are favored to develop various types of VQA reviewed in Ref. [10, 4]. In essence, what makes VQE (generally VQAs) engrossing is the trial state analysis step to choosing an ansatz adaptive to particular quantum device architecture and/or problems. Subsequently, the suitable ansatz is executed congruently on the NISQ device, the remainders are then handled by the classical computer with the hybrid loop to generate remarkable results with the aid of error mitigation techniques. Instead of diagonalizing a matrix representation of Hamiltonian $\hat{H}$ to find out eigenvalues that exponentially scale up the matrix size and calculation overhead in traditional computing as the problem grows up larger, VQE just focuses on catching up with the ground state energy $E_{g}$ basically based on the variational principle
@@
The arbitrary quantum state $|\Psi({\bm{\theta}})\rangle\equiv\hat{U}({\bm{\theta}})|\Psi_{0}\rangle$ is a trial solution, the so-called ansatz parametrized by a unitary operator $\hat{U}({\bm{\theta}})$, so that when the parameter ${\bm{\theta}}$ varies, the ansatz $|\Psi({\bm{\theta}})\rangle$ is readily capable to span on the relevant quantum space where the ground state located. By modifying ${\bm{\theta}}$ value systematically, the VQE task is to travel around spanned space to extract the ground state minimizing the energy function $E\left[\Psi({\bm{\theta}})\right]$ or, in general VQA, the energy function described by $\mathcal{L}\left({\bm{\theta}},\langle\hat{H}\rangle_{\hat{U}({\bm{\theta}})}\right)$, and the objective is
@@
In fact, the optimal wave function is not necessarily required to close to the ground state solution, due to the error in energy being in the second order of the quantum state error. Formally, the classical approach for VQE insists on an analytical expression of the lost function $\mathcal{L}\left({\bm{\theta}},\langle\hat{H}\rangle_{\hat{U}({\bm{\theta}})}\right)$, in particular, the wave function expression $|\Psi(\bm{\theta})\rangle$ that becomes eventually unachievable going along with more and more problem complexities. Nevertheless, this weakness is literally the quantum advantage we have in the VQE quantum computing approach. Quantum computers are adopted to simulate quantum states, and state expectation outcomes are subsequently measured to pipeline to calculate the lost function classically. And partly because of lost function precision greater quadratically than the state function, indeed, we need two states to sandwich operators, VQEs are hereby friendly with noisy devices, which can work out meaningful results even though the states are disturbed by noises. These motivations are robust enough to let us dive deeper into VQE operation, ideally with a noise-free assumption in this paper’s scope though, the noisy one will be addressed in the future study.
@@
The structure of a VQE algorithm is depicted schematically as Fig. 1 within four basic steps: Hamiltonian construction, Ansatz preparation, Measurement strategy and Optimization.
@@
Hamiltonian construction: Starting from a given abstract problem, eg. molecule ground state energy, shortest path, transshipment problems, an initially mathematical form called the Hamiltonian $\hat{H}$ is modeled. A mapping scheme is needed to convert $\hat{H}$ into a new form made out of operators that are able to execute on a quantum computer. Invoking prior knowledge of the physical system, particular symmetries are drawn on to simplify the Hamiltonian, such information is usefully reused to guide the erection of other VQE subroutines.
@@
Ansatz preparation: To pick out a good ansatz, an essential condition that must be satisfied is that the space spanned by the ansatz contains the desired state to extremize the objective function. An obvious one is a generic ansatz spread over all Hilbert space, a quantum circuit for that kind of ansatz is usually generated by multi-control-$U_{3}({\bm{\theta}})$ gates so that all problems could be ultimately settled without ansatz concern. Nevertheless, that is not the end of the story, the computation cost of such a multi-control-$U_{3}({\bm{\theta}})$ gate is extremely high and even unfeasible when deployed on the NISQ device. As a consequence, an initial crucial step in VQE is to find an ansatz that is expressive enough to capture the important features of the wavefunction but also efficient enough to be optimized using optimization techniques. In this article, we employ the approach that excavates up symmetries and physical properties of the physical system to guide the design of the ansatz respecting those symmetries, which helps us considerably reduce the number of training parameters and circuit depth needed in a theoretically speculative way, thereby improves the efficiency of the optimization. There are also other methods that use a hierarchical ansatz, where the circuit is divided into layers and each layer is optimized independently, and also use machine learning to automatically generate the quantum circuit [11, 12].
@@
Measurement: To read off information from the wave function to estimate the expectation value of the Hamiltonian, apart from the standard basis measurement where each qubit in the output state is measured in the computational basis, the measurement scheme is manifestly came up with is a unitary transformation to the diagonal basis of observable operators in the Hamiltonian, typically defined by tensor products of Pauli operators, such as X, Y, and Z. Almost Hamiltonian of interest are hermitian and be able to decompose in terms of Pauli strings
@@
each Pauli string term $P^{i}=\otimes_{j=1}^{N}\sigma^{i}_{j}$ has likely overlapped the others in somewhere qubit sites, which means $\sigma^{m}_{j}\equiv\sigma^{k}_{j}$ being able to measure simultaneously, the grouping methods skillfully make use of this information to lower the number of measurements, the symmetry structures of the problem are able to exploit as well, and other efficient measurement methods are reviewed in [13]. Those on-research topics are highly crucial to the current NISQ device when a minimum measurement performance is required upon a given restricted number of shots. Furthermore, for a given $\epsilon$ accuracy, VQE demands a trade-off between the $\mathcal{O}(\frac{1}{\epsilon^{2}})$ number of circuit samples and $\mathcal{O}(1)$ circuit depth with other finding ground state energy methods such as QPE, which consumes $\mathcal{O}(1)$ circuit measurement repetitions with depth $\mathcal{O}(\frac{1}{\epsilon})$. A better approach is essentially proposed to reconcile disadvantages between the two above methods, which is deemed a generalized VQE algorithm, $\alpha$-VQE, when outrunning VQE by just $\mathcal{O}(\frac{1}{\epsilon^{2(1-\alpha)}})$ circuit samples and lower circuit depth $\mathcal{O}(\frac{1}{\epsilon^{\alpha}})$ than QPE, where $\alpha\in\left[0,1\right]$ [14].
@@
Optimization: Last but not least, the last piece in VQE subroutine is to vary the parameter ${\bm{\theta}}$ to catch up with the global extreme point (but in many cases, the local one is preferred) in the function of interest described by Eq. (2). The criteria should be pondered to problem-friendly implement, which are able to low-cost function evaluation, fast convergence with a given quantum of resources, and silence to noise. Based on different kinds of utilizing objective function information, people classify it into two main optimization approaches: gradient-based and derivative-free. Meanwhile, the gradient-based employs the gradient information in mostly first order, and second order to travel along the steepest direction of the cost function landscape. The derivative-free rather directly uses responses of the objective function at different points in the parameter space to iteratively improve the parameter values, and as usual, these utilize the subroutine classical algorithm and just consume a limited quantum evaluation expense than gradient-based, these are thus seemingly more congruous with NISQ. In the context of VQAs, gradient-free approaches can be used to optimize quantum circuit parameters when computing the gradient is either impractical or excessively costly, for example, when simulating the behavior of a quantum system becomes computationally expensive, or when the objective function cannot be expressed analytically. Nevertheless, the gradient-based methods are arguably more reliable and have faster convergence as the more complicated structure of the objective function, henceforth, the total trade-off would be shoulder-by-shoulder when deliberately adopting between these two. Furthermore, the step size problem in gradient-based and applying machine learning in derivative-free are active topics of research at this stage, see [4] section II.D for more reviews. In the context of this paper, we investigated several optimization methods and instead labeled them again into two regimes, classical and quantum optimization.
III Ansatz for Transverse Ising model
III.1 Transverse Ising model (TIM)

### v2

# Variational Quantum Eigensolver: A Comparative Analysis of Classical and Quantum Optimization Methods

@@ Abstract
In this study, we study the Variational Quantum Eigensolver (VQE) application for the Ising model as a test bed model, in which we pivotally delved into several optimization methods, both classical and quantum, and analyzed the quantum advantage that each of these methods offered, and then we proposed a new combinatorial optimization scheme, deemed as QN-SPSA+PSR which combines calculating approximately Fubini-study metric (QN-SPSA) and the exact evaluation of gradient by Parameter-Shift Rule (PSR). The QN-SPSA+PSR method integrates the QN-SPSA computational efficiency with the precise gradient computation of the PSR, improving both stability and convergence speed while maintaining low computational consumption. Our results provide a new potential quantum supremacy in the VQAs’s optimization subroutine, even in Quantum Machine Learning’s optimization section, and enhance viable paths toward efficient quantum simulations on Noisy Intermediate-Scale Quantum Computing (NISQ) devices. Additionally, we also conducted a detailed study of quantum circuit ansatz structures in order to find the one that would work best with the Ising model and NISQ, in which we utilized the properties of the investigated model.
@@
Ising Model, Variational Quantum Algorithm, Quantum Optimization, Ansazt Construction, Gradient Estimation
@@ I Introduction
From the early stage of studying quantum computers, that by its very quantum nature is promising for a bright new age of computation coming along with three crucial keys of quantum theory: quantum probabilistics, superposition, and entanglement. The distinct properties of quantum computers (QC) from classical computers make it be more unique in application, not only be an upgraded computational speed version of traditional computers. Currently, when the quantum hardware is rather limited on operational stats and noise resilience, despite of the ambiguity of the realization of such advantages, Variational Quantum Algorithms (VQAs) [1, 2] are thought to be the best at outperforming conventional computers, which are able to implement well on near-term quantum devices known as the so-called Noisy Intermediate-Scale Quantum (NISQ) computers [3, 4, 5, 6, 7].
@@
Nowadays, when the quantum revolution is being on track, big tech companies like IBM, Google, D-Wave, etc. compete to build their own quantum computers to yield quantum supremacy, where the first-ever experimental demonstration was achieved by the Google AI Quantum team [8], but that is still far from what we expect the quantum computer could do, while coherence time, connectivity of qubits, and qubit number limitations keep us from being able to successfully run a long-depth circuit and produce a significant result on the present noisy device, in addition to quantum gate implementation issues. Those restrictions would yet be challenging to hardware scientists in the near future, posing obstacles that computer scientists would need to weigh against trainability, precision, and efficiency, nevertheless, these NISQ devices are capable of exploitation, see Ref. [9] for further reviews. Standing out among quantum algorithms envisioned to beat classical computers, which are mostly designed for the fault-tolerant quantum computer, VQAs turn up regarded as an appropriate candidate compatible with the current defective quantum device to address these constraints [2]. The first two prominent applications of VQAs coming up are the Variational Quantum Eigensolver (VQE) [10] and the Quantum Approximate Optimization Algorithm (QAOA) [11]. These inherit the core scheme of a VQA, which is to utilize the hybrid routine, leverage the versatility and computational power of classical computers to handle the computation processing, and use quantum devices to execute the quantum circuit. Hence, in this paper’s scope, we will travel through the standard procedures of VQE and investigate the properties of different optimization methods implemented in the VQE’s subroutine.
@@
The structure of this paper is as follows: in Chapter 2, we provide an overview of the Variational Quantum Eigensolver (VQE) procedures, a brief introduction to the Transverse Ising Model, and its symmetries are also discussed here, which will be utilized to inform the ansatz structure. Subsequently, Chapter 3 delves into optimization methods, encompassing both classical and quantum approaches, where we will propose a new quantum optimization method. The numerical study and experiment are conducted in Chapter 4, and then we make the final conclusion in Chapter 5.
II Variational Quantum Eigensolver for Ising model
II.1 Overview
Figure 1: Variational Quantum Eigensolver (VQE) Architecture. This schematic representation illustrates the workflow of the VQE algorithm, which is a hybrid quantum-classical optimization technique used to estimate the ground state energy of a given Hamiltonian. The process consists of four key steps: Hamiltonian Construction (Step 1): The problem is encoded into a Hamiltonian $\hat{H}$, expressed as a weighted sum of Pauli operator products acting on N qubits; Ansatz Preparation (Step 2): A parametrized quantum circuit is constructed by unitary operator $\hat{U}({\bm{\theta}})$, where $\bm{\theta}$ represents tunable parameters; Measurement Strategy (Step 3): The expectation value $\langle\hat{H}\rangle$ is obtained by applying a basis rotation followed by a specific measurement scheme; Optimization (Step 4): The measured expectation value is fed into a classical optimizer, which minimizes energy $\langle\hat{H}\rangle_{\hat{U}({\bm{\theta}})}$ by updating the parameters $\bm{\theta}$, this step is the crucial point that we will dig into details within this study. The Hybrid Loop, where quantum computations (state preparation and measurement) are performed on a quantum processor, while classical optimization is handled by a classical computer. The updated parameters $\bm{\theta}_{\text{new}}$ are fed back into the quantum circuit iteratively.
@@ II.1 Overview
VQE was initially proposed by Peruzzo et al. [10, 12] as an efficient alternative way to compute the ground state energy, or expanded to estimate the higher-energy levels [13, 14, 15, 16, 17], of many quantum chemistry systems [18, 19] or many-body systems, i.e. H2 [20, 21], LiH and BeH2 [22] molecules, rather than Quantum Phase Estimation (QPE) [23] which demands impractically huge numbers of quantum gates. Besides that, applications of VQE in drug discovery [24, 25] and materials [26, 27, 28] are considerably concerned. Consequently, VQE becomes a highly flexible, viable strategy with current potential quantum resources that can outpace conventional computers [29, 30]. Following the success of running on the photonic quantum processor combined with the traditional devices, many research works favor the development of various types of VQA reviewed in Ref. [31, 9]. In essence, what makes VQE (generally VQAs) engrossing is the trial state analysis step to choosing an ansatz adaptive to particular quantum device architecture and/or problems. Subsequently, the suitable ansatz is executed congruently on the NISQ device, the remainders are then handled by the classical computer with the hybrid loop to generate remarkable results with the aid of error mitigation techniques. Instead of diagonalizing a matrix representation of Hamiltonian $\hat{H}$ to find out eigenvalues that exponentially scale up the matrix size and calculation overhead in traditional computing as the problem grows up larger, VQE just focuses on catching up with the ground state energy $E_{g}$, basically based on the variational principle
@@
The arbitrary quantum state $|\Psi({\bm{\theta}})\rangle\equiv\hat{U}({\bm{\theta}})|\Psi_{0}\rangle$ is a trial solution, the so-called ansatz parametrized by a unitary operator $\hat{U}({\bm{\theta}})$, so that when the parameter ${\bm{\theta}}$ varies, the ansatz $|\Psi({\bm{\theta}})\rangle$ is readily capable of spanning on the relevant quantum space where the ground state is located. By modifying ${\bm{\theta}}$ value systematically, the VQE task is to travel around spanned space to extract the ground state minimizing the energy function $E\left[\Psi({\bm{\theta}})\right]$ or, in general VQA, the energy function described by $\mathcal{L}\left({\bm{\theta}},\langle\hat{H}\rangle_{\hat{U}({\bm{\theta}})}\right)$, and the objective is
@@
In fact, the optimal wave function is not necessarily required to be close to the ground state solution, due to the error in energy being in the second order of the quantum state error. Formally, the classical approach for VQE insists on an analytical expression of the loss function $\mathcal{L}\left({\bm{\theta}},\langle\hat{H}\rangle_{\hat{U}({\bm{\theta}})}\right)$, in particular, the wave function expression $|\Psi(\bm{\theta})\rangle$ that eventually becomes unachievable going along with more and more problem complexities. Nevertheless, this weakness is literally the quantum advantage we have in the VQE quantum computing approach. Quantum computers are adopted to simulate quantum states, and state expectation outcomes are subsequently measured to pipeline to calculate the loss function classically. And partly because of the loss function precision being greater quadratically than the state function, indeed, we need two states to sandwich operators. VQEs are hereby friendly with noisy devices, which can work out meaningful results even though the states are disturbed by noise. These motivations are robust enough to let us dive deeper into VQE operation, ideally with a noise-free assumption in this paper’s scope though, the noisy one will be addressed in a future study.
@@
The structure of a VQE algorithm is depicted schematically as Fig. 1 within four basic steps: Hamiltonian construction, Ansatz preparation, measurement strategy, and Optimization.
@@
Hamiltonian construction: Starting from a given abstract problem, e.g., molecule ground state energy, shortest path, transshipment problems, an initially mathematical form called the Hamiltonian $\hat{H}$ is modeled. A mapping scheme is needed to convert $\hat{H}$ into a new form made out of operators that are able to execute on a quantum computer. Invoking prior knowledge of the physical system, particular symmetries are drawn on to simplify the Hamiltonian, such information is usefully reused to guide the erection of other VQE subroutines.
@@
Ansatz preparation: To pick out a good ansatz, an essential condition that must be satisfied is that the space spanned by the ansatz contains the desired state to extremize the objective function, more details can be found in Ref. [32]. An obvious one is a generic ansatz spread over all Hilbert space, a quantum circuit for that kind of ansatz is usually generated by multi-control-$U_{3}({\bm{\theta}})$ gates so that all problems could be ultimately settled without ansatz concern. Nevertheless, that is not the end of the story, the computation cost of such a multi-control-$U_{3}({\bm{\theta}})$ gate is extremely high and even unfeasible when deployed on the NISQ device. As a consequence, an initial crucial step in VQE is to find an ansatz that is expressive enough to capture the important features of the wavefunction but also efficient enough to be optimized using optimization techniques. In this article, we employ the approach that excavates up symmetries and physical properties of the physical system to guide the design of the ansatz respecting those symmetries, which helps us considerably reduce the number of training parameters and circuit depth needed in a theoretically speculative way, thereby improving the efficiency of the optimization. There are also other methods that use a hierarchical ansatz, where the circuit is divided into layers and each layer is optimized independently, and also use machine learning to automatically generate the quantum circuit [33, 34].
@@
Measurement: To read off information from the wave function to estimate the expectation value of the Hamiltonian, apart from the standard basis measurement where each qubit in the output state is measured in the computational basis, the measurement scheme is manifestly came up with is a unitary transformation to the diagonal basis of observable operators in the Hamiltonian, typically defined by tensor products of Pauli operators, such as X, Y, and Z. Almost all Hamiltonians of interest are Hermitian and able to decompose in terms of Pauli strings.
@@
each Pauli string term $P^{i}=\otimes_{j=1}^{N}\sigma^{i}_{j}$ has likely overlapped the others in somewhere qubit sites, which means $\sigma^{m}_{j}\equiv\sigma^{k}_{j}$ can be measured simultaneously. The grouping methods skillfully make use of this information to lower the number of measurements, the symmetry structures of the problem are able to exploit as well, and other efficient measurement methods are reviewed in [35]. Those on-research topics are highly crucial to the current NISQ device when a minimum measurement performance is required upon a given restricted number of shots. Furthermore, for a given $\epsilon$ accuracy, VQE demands a trade-off between the $\mathcal{O}(\frac{1}{\epsilon^{2}})$ number of circuit samples and $\mathcal{O}(1)$ circuit depth with other finding ground state energy methods such as QPE, which consumes $\mathcal{O}(1)$ circuit measurement repetitions with depth $\mathcal{O}(\frac{1}{\epsilon})$. A better approach is essentially proposed to reconcile disadvantages between the two above methods, which is deemed a generalized VQE algorithm, $\alpha$-VQE, when outrunning VQE by just $\mathcal{O}(\frac{1}{\epsilon^{2(1-\alpha)}})$ circuit samples and lower circuit depth $\mathcal{O}(\frac{1}{\epsilon^{\alpha}})$ than QPE, where $\alpha\in\left[0,1\right]$ [36].
@@
Optimization: Last but not least, the last piece in the VQE subroutine is to vary the parameter ${\bm{\theta}}$ to catch up with the global extreme point (but in many cases, the local one is preferred) in the function of interest described by Eq. (2). The criteria should be pondered for problem-friendly implementation, which are able to perform low-cost function evaluation, fast convergence with a given quantum of resources, and silence to noise. Based on different kinds of utilizing objective function information, people classify it into two main optimization approaches: gradient-based and derivative-free. Meanwhile, the gradient-based approach employs the gradient information in mostly first order, and second order to travel along the steepest direction of the cost function landscape. The derivative-free rather directly uses responses of the objective function at different points in the parameter space to iteratively improve the parameter values, and as usual, these utilize the classical algorithm subroutine and just consume a limited quantum evaluation expense compared to gradient-based, these are thus seemingly more congruous with NISQ. In the context of VQAs, gradient-free approaches can be used to optimize quantum circuit parameters when computing the gradient is either impractical or excessively costly, for example, when simulating the behavior of a quantum system becomes computationally expensive, or when the objective function cannot be expressed analytically. Nevertheless, the gradient-based methods are arguably more reliable and have faster convergence as the more complicated structure of the objective function, henceforth, the total trade-off would be shoulder-by-shoulder when deliberately adopting between these two. Furthermore, the step size problem in gradient-based and applying machine learning in derivative-free are active topics of research at this stage, see [9] section II.D for more reviews. In the context of this paper, we investigated several optimization methods and instead labeled them again into two regimes: classical and quantum optimization.
II.2 Transverse Ising model